OpenStudy (anonymous):

Can anyone verify this? I'm finding the inverse of the function: f(x)=x²+6x-5 SO far I got this as the inverse: ƒ−1(x) = -3±√(x+14) Here is the part i'm having trouble with, verifying the inverse: ƒ−1(f(x))=x for ƒ−1(f(x)), I was able to verify -3+√(x+14) = x: ƒ−1(f(x)) =-3+√((x²+6x-5)+14) =-3+√(x²+6x+9) =-3+√(x+3)² =-3+(x+3) =x However I'm having trouble in verifying -3-√(x+14)=x This is what I am ending up with: ƒ−1(f(x)) =-3-√((x²+6x-5)+14) =-3-√(x²+6x+9) =-3-√(x+3)² =-3-(x+3) =-3-x-3 =-x-6 Can anyone help me please? Much appreciated, thanks

OpenStudy (anonymous):

ƒ−1(f(x)) =-3-√((x²+6x-5)+14) =-3-√(x²+6x+9) =-3-√(x+3)² =-3-√(-1)²(x+3)² =-3-((-1)(x+3)) =-3-(-x-3) =-3+x+3 =x √9=√3²=√(1x3²)=√((-1)²x3²) Since (-1)²=1 I just played with the numbers. I hope this helps.

OpenStudy (anonymous):

hi firstly thanks for helping me, I'm just now looking at what you did there...

OpenStudy (anonymous):

how did you get -1 out of (x+3)? In step 4?

OpenStudy (anonymous):

if you were factoring -1 out of (x+3). wouldn't it become (-x-3)?

OpenStudy (anonymous):

ƒ−1(f(x)) =-3-√(x+3)² =-3-√1*(x+3)² =-3-√(-1)²(x+3)² (-1)²=1

OpenStudy (anonymous):

* is the multiplication sign

OpenStudy (anonymous):

Sorry, I forgot the braces. ƒ−1(f(x)) =-3-√(x+3)² =-3-√(1*(x+3)²) =-3-√((-1)²(x+3)²)

OpenStudy (anonymous):

ahh i see, so you defined 3² to be equal to (-1)²*(3)² and just substituted (-1)² into the square root then?

OpenStudy (anonymous):

Exactly. ;)

OpenStudy (anonymous):

Wow thank you very much, I never would have known that you could substitute something like under the square root! Thanks again for the help.

OpenStudy (anonymous):

like (-1)²*

OpenStudy (anonymous):

You're welcome. :)

OpenStudy (anonymous):

wait am i to delete this question now?

OpenStudy (anonymous):

You can leave it here. No need to delete it. It may be a reference for persons having a similar mathematical problem. You can post another maths question if you want. :)

OpenStudy (anonymous):

alrighty then, thanks.

OpenStudy (anonymous):

:)